OpenStudy (thatonegirl_):

Determine the intervals on which the graph of F (defined on (0, infinity) is concave up.

OpenStudy (thatonegirl_):

\[F(x)=\int\limits_{0}^{x} \frac{ t }{ 9+t^2 }dt\]

OpenStudy (solomonzelman):

Interesting they gave you an integral of a function. Well, we know that the first derivative determines concavity, right?

OpenStudy (thatonegirl_):

isn't it second deriv?

OpenStudy (thatonegirl_):

and would I use the 2nd fundamental thm of calc for this??

OpenStudy (solomonzelman):

yes, yes ...

OpenStudy (thatonegirl_):

Okay thx

OpenStudy (solomonzelman):

The first derivative increases = concave up The first derivative decreases = concave down

OpenStudy (solomonzelman):

to get the first derivative we don't to integrate and differentiate again, do we ?

OpenStudy (thatonegirl_):

nope

OpenStudy (solomonzelman):

so your first derivative is?

OpenStudy (thatonegirl_):

it would be.. \[\frac{ x }{ 9+x^2 }\]

OpenStudy (thatonegirl_):

right?

OpenStudy (solomonzelman):

yes, it is correct!

OpenStudy (thatonegirl_):

Yayy!! :D

OpenStudy (solomonzelman):

Now the second derivative:)

OpenStudy (thatonegirl_):

quotient rule?

OpenStudy (solomonzelman):

Yes, you may. You also may to re-write with negative exponent if you hate to memorize which function is differentiate first.

OpenStudy (thatonegirl_):

oh okay \[\frac{ 9-x^2 }{ (9+x^2)^2 }\]

OpenStudy (solomonzelman):

Yes.

OpenStudy (solomonzelman):

This is negative when ?

OpenStudy (thatonegirl_):

so now I'd find when it's =0 for critical #s right? then do the test thingy on the number line?

OpenStudy (solomonzelman):

Oh, for critical points you set the first derivative =0. Other possible cases, in general. Any x=a such that f(x) is defined and f'(x) is not. Closed interval boundaries (if any),

OpenStudy (thatonegirl_):

Oh! Is the answer (0,3) ?? concave up?

OpenStudy (thatonegirl_):

It's a continuous function right, so there would be no points when it's undefined?

OpenStudy (solomonzelman):

\(\displaystyle 0<\frac{ 9-x^2 }{ 9+x^2 }\) \(\displaystyle 0<\frac{ (3-x)(3+x)}{ 9+x^2 }\) \(\displaystyle 0<\frac{ (3-x)(3+x)}{ \rm positive }\) So, either both parenthesis are negative or both positive. Both negative, x<-3. Both positive, x>3.

OpenStudy (solomonzelman):

So it is concave up whenever \(x{\tiny~}\in {\tiny~~}[\infty,-3){\tiny~}\cup{\tiny~}(3,\infty)\).

OpenStudy (thatonegirl_):

okay thank you again :))

zepdrix (zepdrix):

Are you sure? :o I think it's the piece in the middle.

OpenStudy (thatonegirl_):

oh yeah wouldn't it be the (-3,3) ? because tht part is positive

OpenStudy (solomonzelman):

Yes, you are right.

OpenStudy (solomonzelman):

9 - x^2 > 0 9 > x^2 |x| < 3

OpenStudy (thatonegirl_):

okay so then on the interval (0,inf) the answer could only be (0,3) right?

OpenStudy (solomonzelman):

stupid parenthesis thinking ... should have just done what I did now at first:)

OpenStudy (thatonegirl_):

lol happens to the best of us xD

OpenStudy (solomonzelman):

obviously if |x|\(\ge\)3, then 9-x^2 is NOT POSITIVE.

OpenStudy (solomonzelman):

Nice catch!

OpenStudy (thatonegirl_):

Thanks you guys!! @SolomonZelman @zepdrix

OpenStudy (thatonegirl_):

1 more question :'D

OpenStudy (solomonzelman):

Sure

OpenStudy (thatonegirl_):

I'll post another one